Find the center of mass of the region bounded by y=x^2 and y=6-x.


1
Expert's answer
2021-04-14T05:37:55-0400

the region bounded by the curves "y=x^2" and "y=6-x" has the points of intervals where;

"x^2 =6-x"

or

"x^2 + x-6 =0 ...(i)"

the quadratic equation (i) solves to give;

x=-3 and x=2 ...(ii)

for a point "x\in(-3,2)" say x=0

"y=x^2 \implies y=0 for x=0"

and

"y=6-x \implies y=6 for x=0"

thus the line y=6-x is above the curve "y=x^2 \forall x\in(-3,2)"

we now assume uniform density "\rho," of the region.

equation (ii) serves as the limits of our finite integrals for area and moment about the axis

we have

"m=\rho A"

"=\rho \intop ^2 _{-3}[6-x-x^2 ]dx"

"= \rho [6x- \frac {x^2}2 - \frac {x^3}3]^2 _{-3}"

"=\rho [\frac {22}3 + \frac {27}2]= \frac {125}6 \rho" units of mass.

moment of the region abotu y=axis is

"m_y=\rho \int ^2 _{-3} (x) [ 6-x-x^2 ]dx"

"=\rho \int ^2 _{-3} [6x-x^2 - x^3]dx"

"=\rho [3x^2 -\frac {x^3} 3 - \frac {x^4}4]^2 _{-3}"

"=\rho [\frac {16}3 - \frac {63}4]= (\frac{ -125}{12} )\rho" units of the length times units of the mass

and the moment abotu the x-axis is

"m_x = \frac12 \rho \int ^2 _{-3} [(6-x)^2 - (x^2)^2]dx"

"= \frac12 \rho \int ^2 _{-3} [36-12x+x^2 - x^4 ]dx"

"=\frac12 \rho [36x-6x^2 + \frac {x^3}3 -\frac {x^5}5]^2_{-3}"

"=\frac12 \rho [\frac {664}{15} -\frac{-612}{5}]= \frac{250}3 \rho" units of length times units of mass

the center of mass ("\bar x" ,"\bar y" ) have

"\bar x= m_y / m =\frac {\frac{-125 \rho}{12} } { \frac{125 \rho} 6 } = -1/2"


"\bar y= m_x /m =\frac { \frac {250\rho}3 } {\frac {125 \rho}6} =4"


"\implies (\frac{-1}2, 4)" is the center of mass of the region bounded by the y=x^2 and y=6-x under the assumption that mass density is uniform over the region.